Q= 1.2C, d= 10cm, I= 10A. The direction of v is perpendicular to I. If v=0 m/s, what is the magnetic force acting on charge Q? If v= 13m/s, what is the magnetic force acting on charge Q? what direction is the force on Q?
Answer #1

if v = 0

magnetic force acting on charge F = QvB sin(theta) = 0

-----------------------

if v = 13 m/s

magnetic field at distance d= 10 cm = 0.1 m

B = Mu0*i/(2pid) = 4pix10^-7*10/(2pi*0.1) = 2x10^-5 T

magnetic force acting on charge , F = 1.2*13* 2 x 10^-5 = 3.12 x10^-4 N